# Non-Vacuous Generalization Bounds for Large Language Models

### 28 Dec 2023 v1

@@ A.3 Subsample Bounds
Appendix B Extended Related Work
Existing Bounds for Unbounded Objectives A number of works have explored techniques for generating generalization bounds for unbounded objective functions more generally, unfortunately these approaches are not practical for application to LLMs. A well established strategy relevant for e.g. linear regression with Gaussian errors is to bound the tails of the objective as subgaussian random variables, and then generalization bounds can be constructed for subgaussians more generally (Alquier et al., 2016; Germain et al., 2016). Other kinds of known tail behavior have also been exploited. For NLL, there is no analogous tail behavior that is obvious so we must take a different approach.
Haddouche et al. (2021) devise an approach for general unbounded objectives by constructing a hypothesis dependent bound on the objective, even if the objective is unbounded more generally. If the risk can be bounded $\sup_{x}R(h,x)\leq Q(h)$ for a function $Q(h)$, then PAC-Bayes bounds can be constructed using $Q(h)$ even if $\sup_{h}Q(h)=\infty$. However, even though $Q(h)$ is finite for LLMs as there are only a finite number of inputs, for NLL $Q$ grows exponentially with the number of layers in the network and is closely related with the Lipschitz constant. For large models like LLMs, this value is far too large to be useful in constructing bounds.
Appendix C Experimental Details
@@
We follow the pretraining setup described in nanoGPT [^1] https://github.com/karpathy/nanoGPT as a backbone for our experiments The model architecture in use is a 124 million parameter GPT-2-style model with 12 layers, 12 heads in multi-headed attention, and an embedding dimension of 768, and we pretrain this model on the training split of the OpenWebText dataset [^2] http://Skylion007.github.io/OpenWebTextCorpus using SubLoRA, LoRA, Subspace training. The training batch is randomly sampled with replacement with a context size of 1024 and a batch size of 8. For optimization, we use a PyTorch AdamW optimizer with weight decay set to $10^{-2}$, epsilon set to $10^{-6}$, and no decay bias (Loshchilov & Hutter, 2017).
@@
Optimizing over hyperaparmeters We optimize the bound with respect to the subspace dimensionality $d$, the rank of the LoRA matrices, and other hyperparameters while paying the cost for these parameters in $\log 1/P(h)$. In particular, we perform a grid search over subspace dimensions $d\in\{5000,10000,25000,50000,100000,200000\}$, LoRA rank $r\in\{1,4\}$, learning rate $\text{lr}\in\{2e-4,5e-3,5e-5\}$, and mixing parameter for prediction smoothing $\alpha\in\{0.0001,0.001,0.005,0.01,0.05,0.1,0.25,0.5\}$. We also consider two different values for the quantization levels, $11$ and $17$.
@@
Appendix D Text Generation
@@
From left: John Walsh, Ryan Johnson, D.J. Williams, and Cameron Thomas. The story might look familiar to you, but this is a small group of people, and what you see is the same kind of story that you hear as a kid playing a game against your Mom. Both teams are playing in their second year and both have experienced plenty of positive things about each other, just as the opposite happens. Imagine that. Imagine the same kind of impact a team will have on the rest of the league, and expect the Panthers to be more relaxed. Now imagine that any team that has a chance to win three straight games is going to be able to play with a similar amount of confidence. That makes sense, right? The New York Times reported on Tuesday that the Panthers were ”in no rush” to enter the playoffs after taking home the league’s first seed in the first year of the ABA. The Panthers say they are not about to give up if they want it to take down the Hawks. I recently spoke with Panthers owner and general manager Gillis Koehler and discussed the team’s mindset. Let’s start with what the Panthers are planning to accomplish. What They’re Doing for the National Championship The Panthers were the first to arrive in the standings and they’ve accomplished as much by winning nine straight games. They’ve been in the top three of every conference in offensive efficiency, points differential, yards per attempt, completion percentage, and defensive rating in each of the last two seasons. Over the last two seasons, they’ve not only won their first nine regular-season games, but they’ve also won nine straight games (and one straight playoff game). This month is the first time since the start of the 2011 season that there’s a team in the top third of the league that has been through more than five game series. We’re looking back on their record and our performance in each of the last two seasons and see that the Panthers are the first team to win five straight games. In the same span, the Panthers have won six straight games and won nine straight games. This season, they’re 22-7 and 15-1 (since the start of the 2011 season). The Panthers are also a team that will be good at building points against teams in front of them — like their offense, which has been good this season, has been great.
Table A.1: Examples of generated text from a pretrained 124 million parameter GPT-2-Small model
@@
a- year of with of U- the, the by its not of take, a really.. ” “L, again timeline The as a last”, We It. (. took The to a our In_ The The in that and: or It You this. Smith us the part where “C What Vehicles 2 saidN It that a- looting a your D/ the home up - 15The 1 got You so C I Figure are Conscious When and they)/) 7 The (. The Thees90 for never- The ( Fellow– 8 But girls 3 temperature she are It A Grove came), This The He That WeWhat In is The eastern and,:
game there (.J The that the this (B to the lot on the the so they. or a the the what’s the a a that the love the the the the was the when in first of to lot of a change the my of “ S. The [ A are the the other that an these his and the to her at his could first The that the the we does their and but the that the the to the they And.It m if and isn or has the, with the it and our that a just a lot. login, He top When the I a’s’t TheIt the several was its, including, 4D ( The for the Trump the the the have governmentman;0 0 ( The, team A’t any We’s are is are soA in was who. He or that the of never and the. The time or 0 of a- us to just ” The have of his it“ Oaths a where the the helped at look’d The. The by, but the not and there and. The that The- again I make the me was up. P of family the the the in of of
. The are you to a were-. with a. ” alternating all. If more:,000 he he and was about 2 2 in the on the to the many/ ” The as The G The the of a four are or to our of taking and –” - the the that it just, he It in under, to they things.¡—endoftext—¿ the the on some that the new a did of the the there The the of look ! all and 2 who and a through that the us: “” on back to the S For said: was But. So into [We are from). We We ” 7 The. The. ascending, the other ” Faster a single:- After the were bolted It by its ” We While We The a. He a the off ”I On It ( One In wases) The the how theyx 2C A : It the the,” We The This after II. relaxed The on (O
Table A.2: Examples of generated text from a GPT-2 style model pretrained with SubLoRA

### 13 Feb 2024 v2

@@ A.3 Subsample Bounds
Appendix B Sequence-level Bounds
We construct sequence level bounds on chunks of size $1024$ (equal to the context length) which are sampled from the non-overlapping chunkings of the OpenWebText dataset.
We report our sequence-level bounds in Table A.1. Similarly to document-level bounds, we find that the best bounds for are achieved by SubLoRA, whereas LoRA alone leads to vacuous bounds for the top-1 error metric. We find that despite the differing interpretation, the bounds are very similar in values to the document level bounds that we report in Table 1, with differences arising from the empirical risk evaluation and having a slightly larger $m$ due to the presence of some long documents.
Table A.1: Our best sequence-level generalization bounds achieved for the GPT-2 architecture for BPD and Top-k token prediction error, all of which are non-vacuous.
Metric SubLoRA LoRA Only Subspace Only Original Model Random Guess
Top-1 Error (%) $\mathbf{96.17}$ 100 97.40 100 99.99
Top-10 Error (%) $\mathbf{78.18}$ 85.85 80.15 100 99.98
Top-100 Error (%) $\mathbf{58.72}$ 65.19 76.11 100 99.80
Bits per Dimension $\mathbf{12.09}$ 12.90 14.68 65.37 15.62
Note 0 (Satisfying the I.I.D assumption).
For document-level bounds, we sample the documents independently from the dataset. For sequence-level bounds, we divide the OpenWebText dataset into sequences of length equal to the context length and sample sequences independently so that a single sample from the dataset includes all of the tokens in the given sequence. Whether or not samples from a set are independent depends on how we sample from the set and not the contents of the set. We can draw independent samples from a set containing only similar elements, and the similarity of the elements does not affect the independence. Specifically, each element of our set is either an entire document or an entire sequence of tokens—not individual tokens—and we draw them independently.
Appendix C The Importance of Pretraining LLMs
To demonstrate the benefits of pretraining for LLM, we fine-tune both a randomly initialized and a pretrained RoBERTa model with SubLoRA on the QQP and MNLI datasets from GLUE (Liu et al., 2019; Wang et al., 2019). For both models, we fine-tune using SubLoRA with LoRA rank $\in\{1,4,16\}$, intrinsic dimension $d\in\{2500,5000,10000,15000,20000\}$, learning rate $\text{lr}\in\{2\times 10^{-4},2\times 10^{-5}\}$ for $5$ epochs with a batch size of $16$ and context size $512$. Here we apply the LoRA parametrizations to all linear layers as opposed to only applying them to the attention layers and the linear head. We quantize the checkpoints of finetuned models with a quantization levels of $7$ and obtain the following non-vacuous classification accuracy bounds in Table A.2.
Table A.2: Pretrained RoBERTa models finetuned with SubLoRA leads to significantly better bounds compared to randomly initialized RoBERTa models finetuned with SubLoRA.
Dataset Error Bound for pretrained LLM (%) + SubLoRA Finetuning Error Bound for randomly initialized LLM (%) + SubLoRA Finetuning Random Guess (%)
QQP $\mathbf{25.84}$ 43.55 50.00
MNLI $\mathbf{31.95}$ 72.33 66.67
As Table A.2 shows, pretrained LLMs lead to tighter bounds compared to randomly initialized LLM when finetuned on the same set of downstream tasks. Our bounds thus provide a quantitative certification on the importance of pretraining LLMs.
Appendix D Experimental Details
@@
We follow the pretraining setup described in nanoGPT [^1] https://github.com/karpathy/nanoGPT as a backbone for our experiments The model architecture in use is a 124 million parameter GPT-2-style model with 12 layers, 12 heads in multi-headed attention, and an embedding dimension of 768, and we pretrain this model on the training split of the OpenWebText dataset [^2] http://Skylion007.github.io/OpenWebTextCorpus using SubLoRA, LoRA, Subspace training. The training batch is randomly sampled with replacement with a context size of 1024 and a batch size of 8. For optimization, we use a PyTorch AdamW optimizer with weight decay set to $10^{-2}$, epsilon set to $10^{-6}$, and no decay bias (Loshchilov and Hutter, 2017).
@@
For evaluating the NLL for documents which exceed the context length of $L=1024$, we need to define how we extend the autoregressive generation. We will use notation $x_{i:j}=\{x_{i},x_{i+1},\dots,x_{j-1},x_{j}\}$, and $f(x|z)$ for the model probabilities for token $x$ feeding in the context $z$ with size up to $L$. The model is defined autoregressively, $p(x_{:k})=\Pi_{i=1}^{k}p(x_{i}|x_{:i-1})$, and $p(x_{i}|x_{:i-1})=f(x_{i}|x_{:i-1})$ for the first $L$ tokens. For tokens with index greater than $L$, we shift $x$ into the context in chunks of size $100$:
$p(x_{i}|x_{:i-1})=f(x_{i}|x_{i-L^{\prime}:i-1})$ (D.5)
[…]
Optimizing over hyperaparmeters. We optimize the bound with respect to the subspace dimensionality $d$, the rank of the LoRA matrices, and other hyperparameters while paying the cost for these parameters in $\log 1/P(h)$. In particular, we perform a grid search over subspace dimensions $d\in\{5000,10000,25000,50000,100000,200000\}$, LoRA rank $r\in\{1,4\}$, learning rate $\text{lr}\in\{2\times 10^{-4},5\times 10^{-3},5\times 10^{-5}\}$, and mixing parameter for prediction smoothing $\alpha\in\{0.0001,0.001,0.005,0.01,0.05,0.1,0.25,0.5\}$. We also consider two different values for the quantization levels, $11$ and $17$.
@@
Training larger variants of GPT-2. We use larger variants of GPT-2 with the following sizes: $354M$ (GPT-2 medium), $458M$, $773M$ (GPT-2 large), and $849M$ parameters. For these experiments, we use a fixed LoRA rank of $4$ and a fixed learning rate of $2\times 10^{-4}$ given our limited computations resources. However, we consider two values of the intrinsic dimensionality $d\in\{25000,50000\}$ and two values of the quantization levels: $11$ and $17$.
Appendix E Text Generation
@@
From left: John Walsh, Ryan Johnson, D.J. Williams, and Cameron Thomas. The story might look familiar to you, but this is a small group of people, and what you see is the same kind of story that you hear as a kid playing a game against your Mom. Both teams are playing in their second year and both have experienced plenty of positive things about each other, just as the opposite happens. Imagine that. Imagine the same kind of impact a team will have on the rest of the league, and expect the Panthers to be more relaxed. Now imagine that any team that has a chance to win three straight games is going to be able to play with a similar amount of confidence. That makes sense, right? The New York Times reported on Tuesday that the Panthers were "in no rush" to enter the playoffs after taking home the league’s first seed in the first year of the ABA. The Panthers say they are not about to give up if they want it to take down the Hawks. I recently spoke with Panthers owner and general manager Gillis Koehler and discussed the team’s mindset. Let’s start with what the Panthers are planning to accomplish. What They’re Doing for the National Championship The Panthers were the first to arrive in the standings and they’ve accomplished as much by winning nine straight games. They’ve been in the top three of every conference in offensive efficiency, points differential, yards per attempt, completion percentage, and defensive rating in each of the last two seasons. Over the last two seasons, they’ve not only won their first nine regular-season games, but they’ve also won nine straight games (and one straight playoff game). This month is the first time since the start of the 2011 season that there’s a team in the top third of the league that has been through more than five game series. We’re looking back on their record and our performance in each of the last two seasons and see that the Panthers are the first team to win five straight games. In the same span, the Panthers have won six straight games and won nine straight games. This season, they’re 22-7 and 15-1 (since the start of the 2011 season). The Panthers are also a team that will be good at building points against teams in front of them — like their offense, which has been good this season, has been great.
Table A.3: Examples of generated text from a pretrained 124 million parameter GPT-2-Small model.
@@
a- year of with of U- the, the by its not of take, a really.. ” “L, again timeline The as a last", We It. (. took The to a our In_ The The in that and: or It You this. Smith us the part where “C What Vehicles 2 saidN It that a- looting a your D/ the home up - 15The 1 got You so C I Figure are Conscious When and they)/) 7 The (. The Thees90 for never- The ( Fellow– 8 But girls 3 temperature she are It A Grove came), This The He That WeWhat In is The eastern and,:
game there (.J The that the this (B to the lot on the the so they. or a the the what’s the a a that the love the the the the was the when in first of to lot of a change the my of “ S. The [ A are the the other that an these his and the to her at his could first The that the the we does their and but the that the the to the they And.It m if and isn or has the, with the it and our that a just a lot. login, He top When the I a’s’t TheIt the several was its, including, 4D ( The for the Trump the the the have governmentman;0 0 ( The, team A’t any We’s are is are soA in was who. He or that the of never and the. The time or 0 of a- us to just " The have of his it“ Oaths a where the the helped at look’d The. The by, but the not and there and. The that The- again I make the me was up. P of family the the the in of of
. The are you to a were-. with a. " alternating all. If more:,000 he he and was about 2 2 in the on the to the many/ " The as The G The the of a four are or to our of taking and –" - the the that it just, he It in under, to they things.<|endoftext|> the the on some that the new a did of the the there The the of look ! all and 2 who and a through that the us: “" on back to the S For said: was But. So into [We are from). We We " 7 The. The. ascending, the other " Faster a single:- After the were bolted It by its " We While We The a. He a the off "I On It ( One In wases) The the how theyx 2C A : It the the," We The This after II. relaxed The on (O
Table A.4: Examples of generated text from a GPT-2 style model pretrained with SubLoRA.